A big barrel contains 300 L of a 6% solution of lawn fertilizer. How much water should be added to obtain a 5% solution?
Answer by mananth(16946)

Acid 6 % ----------300 L
water 0% ---------- x L
required 5% ------ Total will be x+300
..
6*300+0*x=5(x+300)
1800 =5x+1500
-1500
1800-1500=5x
300 =5x
/5
300/5 = x
60 = x
60 liters of water
...
m.ananth@hotmail.ca